OpenStudy (artemisxrpg):

Earth revolves on its axis once every 24 hr. Assuming that Earth's radius is 6400 km, find the following. a. angular speed of earth in radians per day and radian per hours b. linear speed at the north pole or south pole c. linear speed at quito,ecuador, a city on the equator d. linear speed at salem, oregon (halfway from the equator to the north pole)

ganeshie8 (ganeshie8):

\(\large \omega = \frac{2 \pi}{T}\)

ganeshie8 (ganeshie8):

earth is taking 1 day to to rotate around itself. \(\large \omega = \frac{2 \pi~radians}{1~ day}\)

OpenStudy (artemisxrpg):

it's 2 pi because that how much one revolution is?

ganeshie8 (ganeshie8):

Yup ! to go around circle once, it takes 2pi radians

ganeshie8 (ganeshie8):



OpenStudy (artemisxrpg):

so angular speed in 24 hrs would be pi/12

ganeshie8 (ganeshie8):

Yes, pi/12 radians per hour